OpenStudy (anonymous):

Sixty percent of a roof, or 1500 square feet of the roof, has been re-shingled. What is the total area of the roof?

OpenStudy (anonymous):

@ShellD93 @random231

OpenStudy (anonymous):

2500 sq ft used ratio and proportion solution for this 60/1500 = 100(whole percent of a roof)/x (what you need to find) 60x= 100(1500) 60x=150000 x=150000/60 x=2500 :)

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

can u help w/ another one??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Solve: 6|y| + 6 = 4 a.) {–6, 4} b.){–4, 4} c.){0} d.) no solution

OpenStudy (anonymous):

sorry idk.. but i think it is d

OpenStudy (anonymous):

thanks